LauraMalcolm001:

Celia simplified the expression Negative 4 x (1 minus 3) minus (2 x + 2). She checked her work by letting x = 5 in both expressions. Her work for the original expression is shown below. Negative 4 x (1 minus 3) minus (2 x + 2) = negative 4 (5) (1 minus 3) minus (2 (5) + 2) = negative 20 (negative 2) minus 12 = 40 minus 12 = 28. Celia knows that the simplified expression should also equal 28 when x = 5. Which shows Celia’s simpllifed expression? –6x – 1 –6x + 2 6x + 1 6x – 2

supie:

Ok so what we have to do is get the equation which would be \[(−4x)(1−3)−(2x+2)\] So first we have to distribute the - sign \[(−4x)(1−3)+−1(2x+2)=\]\[(−4x)(1−3)+−1(2x)+(−1)(2)=\]\[(−4x)(1−3)+−2x+−2=\]\[8x+−2x+−2=\] then we combine the like terms \[8x+−2x+−2=\]\[(8x+−2x)+(−2)=\]\[6x+−2=?\]

LauraMalcolm001:

6x+-2 could also be considered 6x-2 right?

supie:

Yes! Good work.